OpenStudy (anonymous):

The degree measures of one angle of a parallelogram is two times that of another angle. The angles are consecutive. Find the degree measure of each angle of a parallelogram.

OpenStudy (anonymous):

x + 2x = 180 degree 3x = 180 degree x = 180/3 x = 60 so one angle 60 other 2*60 = 120